The inlet pipe to the pump is 0.06 m in diameter commercial steel and 8.5 m in length with a
fanning friction factor of 0.0046. The discharge line consists of 60 m of 0.06 m diameter steel
pipe. All valves are fully open globe valves. Determine the flow rate through the system.
1
Expert's answer
2020-12-22T05:06:06-0500

Hf=f *l/d*v2/(2g)

Low level height=3 m

High level height=10 m

Thus Hf=10-3=7m

7=0.0046*(60+8.5)/0.06*v2/(2*9.81)

7=0.0046*68.5/0.06v2/20

7*0.06*20=0.0046*68.5*v2

v=5.16m/s

Thus Q=A*v=3.14*d2/4*5. 16

=3.14*0.062/4*5.16

=0.0145888401 m3/s